Ariana took 2 hour longer to drive 360 miles on the first day of her trip then she took to drive 250 miles on the second day. If her speed was the same on both days, what was the driving time each day?
360-250/2=110/2=55
is that right?

Answer by edjones(8007)

d=st d=distance t=time s=speed
360=s(t+2)
s=360/(t+2)
.
250=st
s=250/t
.
360/(t+2)=250/t
360t(t+2)/(t+2)=250t(t+2)/t
360t=250(t+2) no more fractions.
360t=250t+500
360t-250t=250t-250t+500
110t=500
110t/110=500/110
t=50/11 hrs
250/(50/11)
=250*11/50
=55 mph
Always check:
55*50/11=250 true
.
Ed